OpenStudy (anonymous):

A new movie theater has stadium seating. The first row has 10 seats. Each additional row has 4 seats more than the row before it. How many seats are in the theater if there are 16 rows of seats?

OpenStudy (anonymous):

This is an arithmetic progression So the first row has 10 The 16th row should have 10+4(15)=70 Total=(16/2)(10+70)=640

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

so i was right :) : 16 x 4 = 64 x 10 = 640 seats

OpenStudy (anonymous):

I don't even know you and your starting with me? -_-